When using the quadratic formula to solve a quadratic equation (ax2 + bx + c = 0), the discriminant is b2 - 4ac. This discriminant can be positive, zero, or negative.
Create three unique equations where the discriminant is positive, zero, or negative. For each case, explain what this value means to the graph of y = ax2 + bx + c.

Answer by LifeDeathnCandy(5)

Discriminant= b^2-4ac
Positive discriminant= x^2+9x+1=0
b^2-4ac = 9^2-4(1)(2)=81-8=72
Meaning= The equation has two solutions because the discriminant is positive,thre are two xcoordinates.
Negetive discriminant= 3x^2+2x+2=0
b^2-4ac = 2^2-4(3)(2)=4-24= -20
Meaning= The equation has no solutions because the discriminant is negetive, there are no xcoordinates.
Zero Discriminant= 4x^2+4x+1=0
b^2-4ac = 4^4-4(4)(1)=16-16=0
Meaning= The equations has 1 solution.
Hope this helps. Good luck.